# PiFlow: Principle-aware Scientific Discovery with Multi-Agent Collaboration

### v1

@@ 7 Conclusion
In conclusion, we propose PiFlow, a Plug-and-Play module to strategically guide the Hypothesis-Validation loop through a principle-steering mechanism to address challenges of aimless hypothesis searching and unclear connections between hypotheses and evidence. Our approach utilizes a Min-Max optimization that explicitly balances exploitation of high-potential principles with exploration of novel hypotheses, guaranteed by a sublinear average regret bound. Extensive experiments demonstrate that PiFlow can adaptively navigate complex hypothesis spaces without premature convergence on suboptimal solutions, yielding significant improvements over baseline methods. In addition, PiFlow enhanced process demonstrates remarkable ability in avoiding local optima, yet keeping structured rationales for better interpretability. We hope PiFlow can contribute to advancing automated scientific discovery, inspiring further exploration and innovation.
Appendix A Comparison of Limitations
While existing methods of LLM-based agent systems demonstrate significant impact in advancing scientific discovery, they limited to direct application in other domains without any changes of architectures and prompts.
@@ Appendix A Comparison of Limitations
The AI Scientist [30] Generating novel research ideas, writes code, executes experiments, visualizes results, describes its findings by writing a full scientific paper AI research
@@
Agent Laboratory [45] Accepting a human-provided research idea and progressing through three stages – literature review, experimentation, and report writing to produce comprehensive research outputs, including a code repository and a research report AI research
@@
CellAgent [52] Constructing LLM-driven biological expert roles - planner, executor, and evaluator - each with specific responsibilities scRNA-seq Analysis
@@
DrugAgent [28] Employing an LLM Planner that formulates high-level ideas and an LLM Instructor that identifies and integrates domain knowledge when implementing those ideas Drug discovery
@@
IAN [35] Leveraging popular pathway and regulatory datasets for protein-protein interactions to perform analysis through a LLM-based multi-agent architecture Protein-protein interactions analysis
@@
Alloy Design AI [14] Automating alloy design and discovery using physics-aware multimodal multi-agent AI Alloy design and discovery
@@
Inorganic Materials AI [47] Accelerating inorganic materials design using generative AI agents. Inorganic materials design
@@
LIDDIA [5] Language-based intelligent agent for drug discovery tasks. Drug discovery
@@
dZiner [3] Rational inverse design of materials facilitated by AI agents. Inverse design of materials
@@
MOOSE-Chem [58] Utilizing Large Language Models for rediscovering unseen chemistry scientific hypotheses. Chemistry hypothesis discovery
@@
OmniScience [39] Domain adaptive pretraining, instruction tuning and reasoning-based knowledge distillation Scientific reasoning
@@
Organic Semiconductor AI [65] Large Language Model-Based AI Agent for research in organic semiconductor devices Organic semiconductor device research
@@
DrugAgent (Explainable) [21] Explainable drug repurposing agent with Large Language Model-based reasoning Drug repurposing
@@
ProtAgents [11] Protein discovery via large language model multi-agent collaborations combining physics and machine learning Protein discovery
To illustrate the challenge of limited generalization, Table 6 outlines several representative LLM-based agent methods. It highlights how their core implementations and prompt engineering strategies are often tightly coupled to specific scenarios, thereby restricting their out-of-the-box applicability to new or diverse problem domains. This tight coupling necessitates significant modifications when attempting to transfer these systems to different contexts, underscoring a key area for future research and development in creating more adaptable and universally applicable AI agents.
Appendix B The Proof of the Convergence
@@
Appendix C Implementation Details
C.1 Algorithm of PiFlow
1
2
Input: $\mathcal{T}_{t}$, and $\lambda_{factor}$
Output: $suggestion$ (string action recommendation)
3
4if $|\mathcal{F}|<3$ then
       return “Initialize one principle to explore. ”
        // Not enough principles
5
6 end if
7
[…]
  // $y_{t}$ in $\mathcal{T}_{t}$ as exploitation score for each $p_{t}$
9 $S_{exploitation}\leftarrow{compute\_exploitation\_scores}(\mathcal{T}_{t})$
  // Embedding differences to others with cosine similarity
10
11for $i\leftarrow 1$ to $|\mathcal{T}|$ do
@@ 9 $S_{exploitation}\leftarrow{compute\_exploitation\_scores}(\mathcal{T}_{t})$
12       $S_{final}[i]\leftarrow(1-\lambda)\cdot S_{exploration}[i]+\lambda\cdot S_{exploitation}[i]$
13
14 end for
15
[…]
17 $p_{best}\leftarrow\mathcal{T}[i_{best}]$
18 $best\_exploitation\_score\leftarrow S_{exploitation}[i_{best}]$
  // $\triangleright$ Score for the chosen principle
19
20if $best\_exploitation\_score>0.7$ then
[…]
@@ 18 $best\_exploitation\_score\leftarrow S_{exploitation}[i_{best}]$
22       $suggestion\leftarrow\text{``Focus on refining: }\frown p_{best}.content\frown\text{High potential.''}$.”
        // $\triangleright$ REFINE
23
24 else if $best\_exploitation\_score>0.4$ then
[…]
26       $suggestion\leftarrow\text{``Validate: }\frown p_{best}.content\frown\text{Moderate potential.''}$.”
        // $\triangleright$ VALIDATE
27
28 else
[…]

### v2

@@ 6 Conclusion
In conclusion, we propose PiFlow, a Plug-and-Play module to strategically guide the Hypothesis-Validation loop through a steering mechanism to address challenges of aimless hypothesizing and unclear connections between hypotheses and evidence. Our approach utilizes a Min-Max optimization that explicitly balances exploitation of high-potential principles with exploration of novel hypotheses, guaranteed by a sublinear average regret bound. Extensive experiments demonstrate that PiFlow can adaptively navigate complex hypothesis spaces without premature convergence on suboptimal solutions, yielding significant improvements over baselines. A detailed discussion of its current limitations and future potential is provided in Appendix C. We hope PiFlow can contribute to advancing automated scientific discovery, inspiring further exploration and innovation.
Appendix A Use of Large Language Models
We utilized a large language model to assist with proofreading and polishing the language in this manuscript.
Appendix B Broader Impact
PiFlow addresses critical bottlenecks in AI-driven scientific discovery where uncertainty leads to aimless exploration. Our method innovatively frames discovery as a structured uncertainty reduction problem. By using an information-theoretic approach within a hypothesis-validation loop, PiFlow systematically filters for instructive scientific principles. Ultimately, PiFlow establishes a new paradigm for automated research, enabling more targeted exploration and accelerating the generation of impactful scientific insights. In Materials Discovery, it speeds the identification of novel compounds like advanced nanomaterials or superconductors, as shown in our tasks. For Biological Discovery, it enhances the search for effective molecules and the understanding of complex systems. Its principles promise similar advancements in other data-intensive fields, from chemistry to medical sciences, facing vast and uncertain hypothesis spaces.
Appendix C Limitation and Future Work
While PiFlow shows notable improvements through its principled Min-Max optimization, its practical implementation approximates a key theoretical component. This means the current system may not fully capture all nuances of true, model-based information gain, especially the direct adversarial interplay with all possible manifestations of the unknown evaluation function $f^{*}$ from the theoretical objective.
Future research could explore more direct estimations of mutual information for this heuristic within the PiFlow framework to potentially further enhance its strategic guidance. Furthermore, we observed that disabling the LLM’s “Thought Mode” surprisingly improves performance, suggesting that forced Chain-of-Thought can induce cognitive fixation. This finding motivates the development of more flexible reasoning frameworks for agents, aiming to better balance deliberate logic with intuitive generation.
Appendix D Detailed Methodological Review
A hallmark and fundamental limitation of many contemporary LLM-based agent systems in scientific discovery is their limited generalizability. As detailed in Table 4, these frameworks are often characterized by a tight coupling between their core logic and a specific scientific domain. This means their implementations, tool integrations, and most critically, their prompt engineering strategies are meticulously tailored for a single scenario, such as organic chemistry or materials science. Although systems like The AI Scientist (Lu et al., 2024) and Agent Laboratory (Schmidgall et al., 2025) demonstrate strong capabilities in scientific research, they are specifically designed for AI domain, along with a focus of the whole workflow rather than strategic decision-making problem. Consequently, transferring these systems to a new domain necessitates significant re-engineering, restricting their out-of-the-box applicability and hindering the development of truly universal systems.
@@ Appendix D Detailed Methodological Review
The AI Scientist (Lu et al., 2024) Generating novel research ideas, writes code, executes experiments, visualizes results, describes its findings by writing a full scientific paper AI research
@@
Agent Laboratory (Schmidgall et al., 2025) Accepting a human-provided research idea and progressing through three stages – literature review, experimentation, and report writing to produce comprehensive research outputs, including a code repository and a research report AI research
@@
CellAgent (Xiao et al., 2024) Constructing LLM-driven biological expert roles - planner, executor, and evaluator - each with specific responsibilities scRNA-seq Analysis
@@
DrugAgent (Liu et al., 2024) Employing an LLM Planner that formulates high-level ideas and an LLM Instructor that identifies and integrates domain knowledge when implementing those ideas Drug discovery
@@
IAN (Nagarajan et al., 2025) Leveraging popular pathway and regulatory datasets for protein-protein interactions to perform analysis through a LLM-based multi-agent architecture Protein-protein interactions analysis
@@
(Ghafarollahi & Buehler, 2025) Automating alloy design and discovery using physics-aware multimodal multi-agent AI Alloy design and discovery
@@
(Takahara et al., 2025) Accelerating inorganic materials design using generative AI agents. Inorganic materials design
@@
LIDDIA (Averly et al., 2025) Language-based intelligent agent for drug discovery tasks. Drug discovery
@@
dZiner (Ansari et al., 2024) Rational inverse design of materials facilitated by AI agents. Inverse design of materials
@@
MOOSE-Chem (Yang et al., 2024c) Utilizing Large Language Models for rediscovering unseen chemistry scientific hypotheses. Chemistry hypothesis discovery
@@
OmniScience (Prabhakar et al., 2025) Domain adaptive pretraining, instruction tuning and reasoning-based knowledge distillation Scientific reasoning
@@
(Zhang et al., 2024b) Large Language Model-Based AI Agent for research in organic semiconductor devices Organic semiconductor device research
@@
DrugAgent (Inoue et al., 2024) Explainable drug repurposing agent with Large Language Model-based reasoning Drug repurposing
@@
ProtAgents (Ghafarollahi & Buehler, 2024b) Protein discovery via large language model multi-agent collaborations combining physics and machine learning Protein discovery
PiFlow (ours) A unified framework for optimizing and searching-oriented scientific discovery problems Board area in science
In contrast, our PiFlow is designed to overcome this challenge by decoupling the strategic decision-making layer from the task execution layer. By architecting PiFlow as a domain-agnostic, Plug-and-Play module, it provides strategic guidance to a minimal hypothesis-testing MAS without embedding domain-specific knowledge within its own logic. This architectural choice yields superior flexibility and obviates the need for extensive, domain-specific prompt engineering for the strategic component, enabling seamless adaptation across diverse scientific fields.
Take-away: PiFlow introduces a domain-agnostic, plug-and-play architecture by decoupling strategic decision-making from domain-specific execution. This design overcomes the critical generalizability limitations inherent in tightly-coupled scientific agent systems.
Appendix E Distinction from Advanced Prompt Engineering
We define a scientific principle following Definition 3.1. Formally, each principle is represented as a structured text proposition that can be algorithmically scored by our MinMax optimization (as exampled in Figure 1). Unlike arbitrary text strings, these principles is supposed to be with logical consistency, and this structural requirement is what distinguishes them from general prompts. In short, principles represent foundational concepts or established patterns within a domain (e.g., “higher hydrophobicity often correlates with better cell membrane penetration”). These “principles”, whether has been validated or not, can be proposed by experts or, crucially, extracted from the LLM’s own vast pre-trained knowledge. They serve as high-level starting points to generate specific, testable hypotheses.
The core distinction of our PiFlow from prompt engineering lies not in the format of the guidance (which is text), but in the algorithmic generation of that guidance:
  • Prompt engineering baseline. The agent quickly became trapped in local optima. Its process was highly repetitive (e.g., repeatedly stating “The system’s behavior is governed by the interplay…”), and it eventually stagnated, making only meaningless tweaks to the g-factor (e.g., 1.1012 $\rightarrow$ $\cdots$ $\rightarrow$ 1.1030). This demonstrates the limitation of a static, unguided hypothesis-testing loop.
  • Systematic breakthroughs with PiFlow. In contrast, the PiFlow-guided agent demonstrated structured, cumulative learning:
    1. 1.
      (Early 1 3 iterations) Principled exploration: It begins with diverse hypotheses to maximize information gain.
    2. 2.
      (Medium stage) Discovery: It identified non-monotonic relationships (e.g., “deviations beyond an optimal configuration reduce chirality”), forming a natural language-based identification.
    3. 3.
      (Late stage) Paradigm Shift: Ultimately, it synthesized the principle of “minimal radius + maximal turns + optimized pitch”, causing a decisive shift in the search space from (fiber_radius=40, helix_radius=70) to (fiber_radius=20, helix_radius=20) and identifying the non-linear sensitivity of the pitch parameter. This unlocked the significant g-factor improvement ($0.84\rightarrow 1.28\rightarrow 1.41\rightarrow 1.51$).
The core of PiFlow is to force the LLM to structure its disorganized internal knowledge into explicit, falsifiable hypotheses. The PiFlow then uses quantitative feedback from experiments to iteratively refine its understanding regarding the task. The value is not in feeding the LLM new knowledge, but in providing a strategic framework to systematically test and organize the knowledge it already possesses, yielding a loop of LLM Knowledge $\rightarrow$ External Evidence $\rightarrow$ Guided LLM Action, rather than LLM $\rightarrow$ LLM.
In summary, the algorithmic core lies in Equation 1 and Algorithm 1: the Min-Max optimization systematically balances regret minimization with information gain maximization. This is operationalized through dynamic principle scoring that updates based on accumulated evidence $\mathcal{T}t={\langle p_{k},y_{k}\rangle}{k=1}^{t}$. Advanced prompt engineering lacks this principled mathematical framework for evidence integration and strategic trade-offs.
Take-away: Instead of relying on static prompts that lead to local optima, PiFlow employs a Min-Max optimization to algorithmically generate and refine structured, falsifiable principles. By systematically integrating experimental feedback, it transforms the LLM’s latent knowledge into a dynamic, self-correcting engine for scientific discovery, enabling cumulative learning and strategic breakthroughs.
Appendix F Proof of the Theorem
@@
Appendix G Algorithmic Realization of the Min-Max Framework
G.1 Practical Proxies for Regret and Information Gain
Algorithm 1 provides a computationally tractable implementation of the abstract Min-Max optimization strategy presented in Eq. 1. It translates the theoretical concepts of regret minimization (exploitation) and information gain maximization (exploration) into concrete, efficiently computable metrics, enabling PiFlow to guide the scientific discovery process effectively. The bridge between theory and practice is established as follows:
Approximating exploitation (minimizing cumulative regret). The theoretical objective of minimizing cumulative regret, $\sum_{t=1}^{T}(v^{*}-f^{*}(h_{t}))$, is centered on favoring principles that consistently yield high-value outcomes $f^{*}(h)$. In our practical implementation, the compute_exploitation_scores function directly approximates this goal. It leverages the historical trajectory of principle-outcome pairs, $\mathcal{T}_{t}=\{\left<p_{k},y_{k}\right>\}_{k=1}^{t}$. The observed outcome $y_{k}$ here serves as a direct proxy for the performance of its corresponding principle $p_{k}$. A principle associated with higher outcomes is considered to have lower regret. Therefore, the resulting normalized score vector ranging from 0 to 1, $S_{exploitation}\in\mathbb{R}^{|\mathcal{T}_{t}|}$, quantifies the empirical success of each principle, and maximizing this score is equivalent to minimizing the cumulative regret based on past evidence.
Approximating exploration (maximizing information gain). The second term in our objective, maximizing the mutual information $I(h_{t};f^{*}|H_{t-1})$, encourages the selection of hypotheses that are most informative about the underlying scientific landscape. A direct computation of mutual information is often intractable. Consequently, Algorithm 1 employs a practical proxy via the compute_exploration_scores function. This is achieved by analyzing the semantic diversity of principles using their sentence level embeddings (e.g., obtained from QwenMax model). We posit that principles which are semantically distant from those already tested are more likely to reveal novel information about the function $f^{*}$. Therefore, we use the cosine distance between a candidate principle’s embedding and the embeddings of previously explored principles as a heuristic for information gain. High exploration scores, $S_{exploration}\in\mathbb{R}^{|\mathcal{T}_{t}|}$, are assigned to conceptually novel principles. This heuristic is theoretically grounded, as detailed in Appendix G.2.
Algorithm 1 Algorithm of PiFlow.
1:Input: $\mathcal{T}_{t}$, and $\lambda_{factor}$
2:Output: suggestionsuggestion (strategic action recommendation)
3:if $|\mathcal{T}_{t}|<3$ then
4:  suggestionsuggestion $\leftarrow$ “Initialize one principle to explore.”
5:else
6:  $S_{\text{exploration}}\leftarrow\text{compute\_exploration\_scores}(\mathcal{T}_{t})$
7:  $S_{\text{exploitation}}\leftarrow\text{compute\_exploitation\_scores}(\mathcal{T}_{t})$
8:  for $i\leftarrow 1$ to $|\mathcal{T}_{t}|$ do
@@ G.1 Practical Proxies for Regret and Information Gain
9:   $S_{\text{final}}[i]\leftarrow(1-\lambda_{factor})\cdot S_{\text{exploration}}[i]+\lambda_{factor}\cdot S_{\text{exploitation}}[i]$
10:  end for
11:  $i_{\text{best}}\leftarrow\arg\max_{i}(S_{\text{final}}[i])$
12:  $p_{\text{best}}\leftarrow\mathcal{T}[i_{\text{best}}]$
13:  $best\_exploitation\_score\leftarrow S_{\text{exploitation}}[i_{\text{best}}]$
14:  if best_exploitation_score>0.7best\_exploitation\_score>0.7 then
15:   $action\_type\leftarrow\text{``refine''}$
@@
16:   $suggestion\leftarrow\text{``Focus on refining: }\frown p_{\text{best}}.content$
17:  else if best_exploitation_score>0.4best\_exploitation\_score>0.4 then
18:   $action\_type\leftarrow\text{``validate''}$
19:   $suggestion\leftarrow\text{``Validate: }\frown p_{\text{best}}.content$
20:  else
21:   $action\_type\leftarrow\text{``explore''}$